I have an astronomy program that provides information on known comets. They provide only two pieces of information: Perhelion and eccentricity.
I have gone through a number of books and web pages both for astronomy problems and math problems trying to find out how to find out the length of the major axis.
I am not very good at working with advanced math (though I have studied some that I needed to use). Please help me with this. Thank you

Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

I have an astronomy program that provides information on known comets. They provide only two pieces of information: Perhelion and eccentricity.
I have gone through a number of books and web pages both for astronomy problems and math problems trying to find out how to find out the length of the major axis.
----
Draw an ellipse with major axis = 2a
---------------------------------------
Put in the center, the focus, and the object at perigee.
If the focus is the sun:
perihelion = distance from focus to object = p
distance of focus from center = c
Eccentricity = c/a
======
a = p + c
a = p + ae
a-ae = p
a(1-e) = p
a = p/(1-e)
2a = 2p/(1-e)
----
You are given p and e so you can determine major axis = 2a
===================
Cheers,
Stan H.
=============

Answer by Alan3354(69443)

I have an astronomy program that provides information on known comets. They provide only two pieces of information: Perhelion and eccentricity.
I have gone through a number of books and web pages both for astronomy problems and math problems trying to find out how to find out the length of the major axis.
------------------
The perihelion is the distance from one of the foci to the intersection of the ellipse its major axis.
The distance from the center to the focus is a*e, where a = 1/2 the major axis and e is the eccentricity.
--------
The perihelion, p = the distance from the focus to a, p = a - f
----------------
f = a*e
p = a-f --> a = p+f
a = p + a*e
a - a*e = p
a(1 - e) = p
a = p/(1 - e)
-------------
major axis = 2p/(1 - e)
That what you're looking for?
-----------------------------
PS Did you see the comet whose perihelion was only 120000 km from the Sun's surface?
Comet Lovejoy.